OpenStudy (anonymous):

HELLPPP! What is the solution of log3x − 5 16 = 2? x = −4 x = 4 x = −3 x = 3

OpenStudy (anonymous):

cant make it out is it \[\huge \log_{3x-5}(16)=2\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok i guess you are supposed to figure out that since \(4^2=16\) you have \[3x-5=4\] and you can solve that for \(x\) in two steps